# Fractional Dynamics in Galactic Nuclei Non-Local Transport, Transient Phenomena and the Nullification of the Schwarzschild Barrier

### v1

@@ Abstract
We demonstrate the necessity of fractional calculus for modeling anomalous diffusion in stellar dynamics, specifically Resonant Relaxation (RR) near a supermassive black hole (SMBH). Standard transport theories rely on the local Fokker-Planck (FP) equation, a formalism restricted to Gaussian processes governed by the Central Limit Theorem (CLT). This approach fails for RR, which is a superdiffusive Lévy flight characterized by non-local transport and governed by the Generalized Central Limit Theorem (GCLT). We apply the space-fractional Fokker-Planck equation (FFPE), utilizing non-local Riemann-Liouville operators, to resolve major discrepancies between local theory and observations. In transient regimes, the FFPE predicts an immediate, linear flux ($\Gamma(t)\propto t$), consistent with high Tidal Disruption Event (TDE) rates in post-starburst galaxies, whereas local FP models predict an unphysical exponential delay ($\Gamma(t)\propto\exp(-C/t)$). Furthermore, we prove mathematically that the non-local integral operators permit “barrier jumping”. This mechanism nullifies transport bottlenecks like the proposed Schwarzschild Barrier (SB). Local models incorrectly interpret the SB as severely suppressing Extreme Mass-Ratio Inspiral (EMRI) rates. We validate our theoretical findings using direct-summation $N$-body simulations incorporating relativistic corrections. The simulations confirm that particle flux remains continuous across the SB in the regime where RR dominates, demonstrating that the barrier does not inhibit EMRI production. The fractional framework provides the correct physical description for non-local transport, resolving long-standing problems concerning TDE and EMRI rates.
@@ 1 Introduction
The statistical nature of RR deviates fundamentally from classical two-body relaxation (subscript “rlx” from here on), which is typically characterized by Brownian motion. RR, in contrast, constitutes a Lévy flight, a random walk characterized by a trajectory of many short steps interspersed with occasional, large jumps whose step-length probability distribution has a heavy, power-law tail.
The distribution of angular momentum step sizes exhibits a heavy power-law tail (see Eq. 2). This behavior signifies a superdiffusive process, where the mean squared displacement (MSD) grows faster than linearly with time,
@@
Significant advancements were made in understanding how relativistic precession affects RR. Direct-summation NN-body simulations heuristically suggested a suppression of RR diffusion at low angular momentum, dubbed the “Schwarzschild Barrier” (SB, D. Merritt et al., 2011). This phenomenon was subsequently addressed by B. Bar-Or & T. Alexander (2014) in terms of Adiabatic Invariance (AI): As put forward by the authors, when the precession period is shorter than the RR coherence time, the coherent torques are averaged out. They developed the $\bm{\eta}$-formalism to calculate effective diffusion coefficients (DCs) by modeling the background potential as correlated noise, leading to the conclusion that the SB severely inhibits EMRI rates (B. Bar-Or & T. Alexander, 2016).
Whilst these approaches provided foundational insights into the interplay between precession and RR, we note that they rely on the assumption that the background potential fluctuations are Gaussian. As explicitly stated in Section 2.2 of B. Bar-Or & T. Alexander (2014), they invoke the CLT to describe the superposition of forces from $N\gg 1$ stars as time-dependent Gaussian random variables. This assumption inherently restricts the analysis to systems with finite variance in the stochastic steps, resulting in a local FP description.
@@
This suppression of high-probability large jumps leads to significant discrepancies in predicting event rates. Local models predict a bottleneck at the SB, suggesting transport across the barrier relies solely on slower two-body relaxation. Furthermore, in transient regimes, such as the refilling of the loss cone (J. Frank & M. J. Rees, 1976) in post-starburst (E+A) galaxies, observations indicate high TDE rates (I. Arcavi et al., 2014; K. D. French et al., 2016; J. Law-Smith et al., 2017), implying rapid replenishment. Local models predict a significant delay, as their inherent step-by-step diffusive nature cannot reproduce this rapid response.
@@
We apply this framework to the aforementioned astrophysical problems. We demonstrate that the non-local nature of the fractional operators resolves the discrepancies inherent in local models. In the transient regime, the fractional framework predicts an immediate, linear increase in the TDE flux. Furthermore, we prove mathematically that the non-local dynamics allow stars to bypass the region of suppressed diffusion via a “barrier jumping” mechanism. We corroborate this theoretical finding with direct-summation $N$-body simulations incorporating post-Newtonian corrections. These simulations explicitly show efficient transport across the SB in the regime dominated by RR. This combined theoretical and numerical evidence demonstrates that the SB does not constitute an actual barrier to RR-driven transport, removing a major theoretical bottleneck for EMRI production.
@@ 2 Anomalous Diffusion and the Fractional Formalism
We analyze the regimes of anomalous diffusion based on the exponent $\alpha$ in Eq. (1). Subdiffusion, corresponding to $\alpha<1$, arises from processes with a heavy-tailed waiting time distribution $\psi(t)\sim t^{-1-\beta}$ (where $0<\beta<1$). This leads to a diverging mean waiting time, effectively ”trapping” particles. This regime is non-Markovian, exhibiting memory effects as discussed in Section 2.3, and could represent stars temporarily caught in resonances. Normal diffusion, $\alpha=1$, represents the classical limit where the mean squared displacement grows linearly with time. This process has finite variance in both jump lengths and waiting times and is governed by the standard Central Limit Theorem. In our context, this corresponds to classical two-body relaxation over long timescales. Finally, superdiffusion, $1<\alpha<2$, arises when the jump length distribution $\lambda(X)$ is heavy-tailed (Eq. 2). This process, known as a Lévy flight, is characterized by an infinite variance in jump lengths. The MSD grows faster than linearly with time. We identify this superdiffusive regime as the correct description for Resonant Relaxation (RR) in angular momentum space.
In RR, coherent torques persist for the precession timescale $T_{\text{precess}}$, generating a Lévy flight characterized by a probability density function (PDF) $P(\Delta J)$ with a power-law tail,
@@ 3 Timescales and Steady-State Implications
We analyze the steady-state regime, where the distribution function $f(J)$ is time-invariant ($\partial f/\partial t=0$). This regime applies to systems in long-term equilibrium or in regions of phase space far from boundaries or barriers, where transport is not limited by the non-local effects discussed in Sections 5 and 6. In this limit, the fractional framework must recover the known results of classical Resonant Relaxation (RR) theory, which serves as a consistency check for the model.
@@ 4.1 The Nature of the Stochastic Process
The limitation stems from the characterization of the stochastic process. Eq. (12) fundamentally describes Brownian motion or generalized Brownian motion (if noise is correlated but Gaussian), which relies on the Central Limit Theorem (CLT) and requires stochastic steps with finite variance. The physical description of RR adopted here corresponds to a Lévy flight (Eq. 2), a process where the variance of the step size is infinite. The evolution of such a system converges to a Lévy stable distribution (Generalized Central Limit Theorem, GCLT), rather than a Gaussian distribution.
While treatments such as B. Bar-Or & T. Alexander (2014) correctly handle the temporal correlations of the noise, they explicitly assume the noise amplitude is Gaussian. In their formalism (Section 2.2 of B. Bar-Or & T. Alexander, 2014), the background potential terms ($h_{nm}^{l}$) are described as time-dependent Gaussian random variables by invoking the CLT, justified by the superposition of forces from $N\gg 1$ stars. This assumption inherently maintains finite variance for the resulting stochastic process and necessitates a description via a local FP equation. This contrasts with the Lévy flight model where the GCLT applies.
Furthermore, one may argue that a pure power-law jump distribution (Eq. 2) is unphysical, as it allows for arbitrarily large jumps. A more physical model might involve “tempered” Lévy flights, where the power law is suppressed by an exponential factor, $P(\Delta J)\sim|\Delta J|^{-1-q}e^{-\lambda J}$ (M. M. Meerschaert & A. Sikorskii, 2011). This process, which converges to a “tempered stable” distribution, possesses finite variance. However, it is still fundamentally non-local and is described by a tempered space-fractional equation utilizing a tempered fractional operator $\mathbb{D}_{x}^{\alpha,\lambda}$ (M. M. Meerschaert & A. Sikorskii, 2011). Crucially, this operator remains an integro-differential operator, maintaining the non-local structure required to model the jumps. The local Fokker-Planck equation, which imposes a Gaussian propagator, is therefore not the correct description even for this more physically realistic, non-local process. This reinforces our central claim: a local differential operator is a fundamentally incorrect choice for describing Lévy-like dynamics, whether pure or tempered.
@@ 6.1 The Schwarzschild Barrier Hypothesis in the Adiabatic Invariance and Local Suppression
Classical two-body relaxation is driven by stochastic encounters and does not depend on orbital coherence. It is entirely unaffected by the SB. However, two-body relaxation is much slower than RR.
The mechanism for RR suppression is understood as Adiabatic Invariance (AI). As detailed by B. Bar-Or & T. Alexander (2014), when the precession frequency exceeds the maximal variability frequency of the (assumed smooth and Gaussian) background noise, the effective diffusion coefficient is exponentially suppressed. This suppression was observed in N-body simulations using Post-Newtonian approximations (D. Merritt et al., 2011) and interpreted as a barrier. The standard view, based on these local models, is that the SB creates a significant bottleneck, as transport across the barrier region is limited primarily to the slow rate of two-body relaxation (B. Bar-Or & T. Alexander, 2016).

### v2

@@ Abstract
We investigate the application of fractional calculus to model stellar dynamics, focusing on Resonant Relaxation (RR) near a supermassive black hole (SMBH). Standard theories use the local Fokker-Planck (FP) equation, restricted to Gaussian processes under the Central Limit Theorem (CLT). We argue this is inadequate for RR. We demonstrate that gravitational interactions inherently produce infinite variance in stochastic torques, violating the CLT. Consequently, RR is governed by the Generalized Central Limit Theorem (GCLT) and constitutes a superdiffusive Lévy flight. We apply the space-fractional Fokker-Planck equation (FFPE), utilizing non-local operators, to explore resolutions to observational discrepancies. In transient regimes, the FFPE predicts immediate, linear flux ($\Gamma(t)\propto t$), consistent with high Tidal Disruption Event (TDE) rates in post-starburst galaxies, whereas local FP models predict significant exponential delay. Furthermore, we demonstrate analytically that non-local integral operators permit “barrier jumping,” bypassing bottlenecks like the Schwarzschild Barrier (SB), which local models interpret as severely suppressing Extreme Mass-Ratio Inspiral (EMRI) rates. We present proof-of-concept $N$-body simulations that confirm non-local RR transport, although the resolution must be improved to rule out enhanced Two-Body Relaxation in the small-N setup. The fractional framework offers a compelling alternative description for non-local transport, potentially resolving TDE and EMRI rate questions.
@@ 1 Introduction
The theoretical understanding of RR, both scalar (changes in magnitude of angular momentum) and vector (changes in orientation, VRR), has advanced significantly through the application of kinetic theory and secular approximations. These approaches often utilize the Balescu-Lenard formalism or related kinetic equations to describe the long-term evolution driven by resonant encounters and correlated noise from potential fluctuations (J.-B. Fouvry & B. Bar-Or, 2018; B. Bar-Or & J.-B. Fouvry, 2018; J.-B. Fouvry et al., 2019a). This framework has enabled the calculation of diffusion coefficients from first principles (B. Bar-Or & J.-B. Fouvry, 2018) and has been applied to complex phenomena such as the coupling of VRR with Lidov-Kozai dynamics (A. S. Hamers et al., 2018). Furthermore, sophisticated numerical methods based on secular dynamics and multipole expansions have been developed to study these processes efficiently (B. Bar-Or et al., 2021). While much of this work focuses on systems amenable to standard kinetic treatments, specific regimes, such as 1D homogeneous systems or certain frequency profiles, exhibit phenomena like kinetic blocking where standard relaxation is suppressed (J.-B. Fouvry et al., 2019b, c).
The statistical nature of RR deviates fundamentally from classical two-body relaxation (subscript “rlx” from here on), which is typically characterized by Brownian motion. We propose that RR exhibits characteristics consistent with a Lévy flight. This type of random walk is characterized by trajectories featuring many short steps interspersed with occasional, large jumps, resulting from a step-length probability distribution with a heavy, power-law tail. We provide a formal theoretical justification for this behavior in Section 2.
If the distribution of angular momentum step sizes indeed exhibits a heavy power-law tail (see Eq. 8), This behavior signifies a superdiffusive process, where the mean squared displacement (MSD) grows faster than linearly with time,
@@
Significant advancements were made in understanding how relativistic precession affects RR. Direct-summation NN-body simulations heuristically suggested a suppression of RR diffusion at low angular momentum, dubbed the “Schwarzschild Barrier” (SB, D. Merritt et al., 2011). This effect was also observed in subsequent numerical studies (e.g., A. S. Hamers et al., 2014). This phenomenon was subsequently addressed by B. Bar-Or & T. Alexander (2014) in terms of Adiabatic Invariance (AI): As put forward by the authors, when the precession period is shorter than the RR coherence time, the coherent torques are averaged out. They developed the $\bm{\eta}$-formalism to calculate effective diffusion coefficients (DCs) by modeling the background potential as correlated noise, leading to the conclusion that the SB severely inhibits EMRI rates (B. Bar-Or & T. Alexander, 2016).
Whilst these approaches provided foundational insights into the interplay between precession and RR, we note that they rely on the assumption that the background potential fluctuations are Gaussian. As explicitly stated in Section 2.2 of B. Bar-Or & T. Alexander (2014), they invoke the CLT to describe the superposition of forces from $N\gg 1$ stars as time-dependent Gaussian random variables. This approach is standard, provided the variance of the individual stochastic contributions is finite. However, as argued in Section 2, the variance of gravitational torques may be infinite. In such cases, the standard CLT does not apply, even for $N\gg 1$. This assumption inherently restricts the analysis to systems with finite variance in the stochastic steps, resulting in a local FP description.
Furthermore, subsequent theoretical work based on kinetic theory, such as B. Bar-Or & J.-B. Fouvry (2018), also adopted local formalisms. These models successfully reproduced the eccentricity diffusion coefficients measured in some numerical simulations, such as those analyzed in A. S. Hamers et al. (2014). This agreement between local theory and simulations warrants discussion. If RR is indeed a Lévy flight, the agreement might indicate that the local models capture the average behavior well in the specific regimes studied, but may fail in scenarios sensitive to the tails of the distribution, such as transport across barriers or transient phenomena.
@@
This suppression of high-probability large jumps leads to significant discrepancies in predicting event rates. Local models predict a bottleneck at the SB, suggesting transport across the barrier relies solely on slower two-body relaxation. This interpretation is based on the observation of the SB in numerical simulations (D. Merritt et al., 2011; A. S. Hamers et al., 2014). We revisit the interpretation of these findings in light of the non-local framework in Section 7. Furthermore, in transient regimes, such as the refilling of the loss cone (J. Frank & M. J. Rees, 1976) in post-starburst (E+A) galaxies, observations indicate high TDE rates (I. Arcavi et al., 2014; K. D. French et al., 2016; J. Law-Smith et al., 2017), implying rapid replenishment. Local models predict a significant delay, as their inherent step-by-step diffusive nature cannot reproduce this rapid response.
@@
We apply this framework to the aforementioned astrophysical problems. We demonstrate that the non-local nature of the fractional operators can potentially resolve the discrepancies inherent in local models. In the transient regime, the fractional framework predicts an immediate, linear increase in the TDE flux. Furthermore, we show analytically that the non-local dynamics allow stars to bypass the region of suppressed diffusion via a “barrier jumping” mechanism. We corroborate this theoretical finding with direct-summation $N$-body simulations incorporating post-Newtonian corrections. These simulations suggest efficient transport across the SB. This combined theoretical and numerical evidence suggests that the SB does not constitute an actual barrier to RR-driven transport, potentially alleviating a major theoretical bottleneck for EMRI production.
2 Lévy Flights and RR
The application of a fractional diffusion framework to Resonant Relaxation (RR) is rooted in the statistical mechanics of the underlying gravitational interactions. In this section we elaborate this argument and demonstrate that the standard assumptions underpinning the local Fokker-Planck formalism are violated by the $1/r^{2}$ nature of gravity, necessitating a description based on Lévy flights and the Generalized Central Limit Theorem (GCLT).
2.1 The Breakdown of the Central Limit Theorem in Gravity
The standard Fokker-Planck equation describes Markov processes where the stochastic fluctuations are Gaussian. This description relies on the Central Limit Theorem (CLT), which states that the sum of many independent and identically distributed (IID) random variables converges to a Gaussian distribution, provided the variance ($\sigma^{2}$) of the individual variables is finite.
However, the gravitational field in a stellar system exhibits fundamentally different statistics. The foundational analysis by J. Holtsmark (1919) and S. Chandrasekhar (1943) established that the probability distribution of the gravitational force $\vec{F}$ is not Gaussian. This is formally derived by analyzing the characteristic function (the Fourier transform of the PDF), $\Phi(\vec{k})=\langle e^{i\vec{k}\cdot\vec{F}}\rangle$. For a 3D isotropic distribution of stars (density $n$), the characteristic function is:
$\Phi(\vec{k})=\exp(-Cn|\vec{k}|^{3/2}),$ (2)
where $C$ is a constant. This is the Holtsmark distribution, a specific type of Lévy $\alpha$-stable distribution with stability index $\alpha=3/2$. The non-analytic dependence on $|\vec{k}|$ (specifically, the exponent $3/2<2$) is a direct consequence of the $1/r^{2}$ force law, which allows the total force to be dominated by rare, strong interactions with nearby stars. The resulting PDF possesses a heavy power-law tail:
$P(F)\sim F^{-5/2},\quad(F\to\infty).$ (3)
2.2 Infinite Variance of Gravitational Torques
We now demonstrate that the torques governing RR inherit this heavy-tailed behavior, leading to a divergence of the variance. We analyze the distribution of torques $P(\tau)$ using the nearest-neighbor approximation, which dominates the tail of the distribution.
In a 3D system, the probability density for the nearest neighbor distance $R$ scales, for small $R$, as $P(R)\propto R^{2}$. The torque magnitude scales as $\tau\propto 1/R^{2}$. We perform a change of variables from $R$ to $\tau$. Let $\tau=C^{\prime}/R^{2}$. Then $R\propto\tau^{-1/2}$, and the Jacobian is $|dR/d\tau|\propto\tau^{-3/2}$. The PDF of the torque is $P(\tau)=P(R(\tau))|dR/d\tau|$:
$P(\tau)\propto R^{2}\,\tau^{-3/2}\propto(\tau^{-1/2})^{2}\,\tau^{-3/2}=\tau^{-5/2}.$ (4)
The distribution of gravitational torques exhibits a heavy tail $P(\tau)\sim\tau^{-(1+\alpha)}$ with index $\alpha=3/2$. We now evaluate the second moment (variance):
$\langle\tau^{2}\rangle=\int_{0}^{\infty}\tau^{2}P(\tau)d\tau\propto\int^{\infty}\tau^{2}\,\tau^{-5/2}d\tau=\int^{\infty}\tau^{-1/2}d\tau.$ (5)
This integral diverges. The variance of the gravitational torque is infinite. More generally, for a Lévy stable distribution with index $\alpha$, all fractional moments $\langle|\tau|^{\theta}\rangle$ diverge if $\theta\geq\alpha$.
2.3 The Generalized Central Limit Theorem and Fractional Kinetics
The divergence of the variance fundamentally invalidates the standard CLT. Consequently, the assumption that the superposition of torques results in a Gaussian process (as invoked in, e.g., B. Bar-Or & T. Alexander 2014 based on $N\gg 1$) is mathematically inconsistent with the statistics of the gravitational field.
When the variance is infinite, the appropriate statistical framework is the Generalized Central Limit Theorem (GCLT) (R. Metzler & J. Klafter, 2000).
GCLT Statement: The sum of IID random variables with a power-law tail $P(X)\sim|X|^{-(1+\alpha)}$, where $0<\alpha<2$, converges to a Lévy $\alpha$-stable distribution.
In RR, the angular momentum changes $\Delta J$ accumulate over time. Assuming the coherence time is not strongly anti-correlated with the torque magnitude, the distribution of $\Delta J$ inherits the heavy tail $P(\Delta J)\sim|\Delta J|^{-5/2}$. Since $\alpha=3/2<2$, the GCLT applies. The resulting stochastic process is a Lévy flight, characterized by superdiffusion and scale invariance.
The kinetic equation governing the evolution of the PDF $f(J,t)$ for a Lévy flight is fundamentally different from the standard Fokker-Planck equation (which corresponds to $\alpha=2$). A Markovian process converging to a Lévy stable distribution must be governed by the space-Fractional Fokker-Planck Equation (FFPE):
$\frac{\partial f(J,t)}{\partial t}=\gamma\frac{\partial^{\alpha}f(J,t)}{\partial|J|^{\alpha}}.$ (6)
The spatial operator is the Riesz fractional derivative, which is a non-local integro-differential operator. Its non-locality is essential to capture the long-range jumps of the Lévy flight. The connection to the underlying statistics is evident in the Fourier domain. The Fourier transform of the Riesz derivative is:
$\mathcal{F}_{k}\left[\frac{\partial^{\alpha}f(J,t)}{\partial|J|^{\alpha}}\right]=-|k|^{\alpha}\mathcal{F}_{k}[f(J,t)].$ (7)
The symbol $|k|^{\alpha}$ is the generator of the Lévy stable process. This structure, specifically with $\alpha=3/2$, directly corresponds to the characteristic function derived from the gravitational interaction (Eq. 2).
@@ 3 Anomalous Diffusion and the Fractional Formalism
We analyze the regimes of anomalous diffusion based on the exponent $\alpha$ in Eq. (1). Subdiffusion, corresponding to $\alpha<1$, arises from processes with a heavy-tailed waiting time distribution $\psi(t)\sim t^{-1-\beta}$ (where $0<\beta<1$). This leads to a diverging mean waiting time, effectively ”trapping” particles. This regime is non-Markovian, exhibiting memory effects as discussed in Section 2.3, and could represent stars temporarily caught in resonances. Normal diffusion, $\alpha=1$, represents the classical limit where the mean squared displacement grows linearly with time. This process has finite variance in both jump lengths and waiting times and is governed by the standard Central Limit Theorem. In our context, this corresponds to classical two-body relaxation over long timescales. Finally, superdiffusion, $1<\alpha<2$, arises when the jump length distribution $\lambda(X)$ is heavy-tailed (Eq. 8). This process, known as a Lévy flight, is characterized by an infinite variance in jump lengths. The MSD grows faster than linearly with time. Based on the arguments presented in Section 2, we identify this superdiffusive regime as the appropriate description for Resonant Relaxation (RR) in angular momentum space.
The coherent torques in RR, persisting for the precession timescale $T_{\text{precess}}$, generate a Lévy flight characterized by a probability density function (PDF) $P(\Delta J)$ with a power-law tail,
@@
(Where $q$ corresponds to the Lévy index $\alpha$ discussed in Section 2, theoretically $q=3/2$.)
@@ 4 Timescales and Steady-State Implications
We analyze the steady-state regime, where the distribution function $f(J)$ is time-invariant ($\partial f/\partial t=0$). This regime applies to systems in long-term equilibrium or in regions of phase space far from boundaries or barriers, where transport is not limited by the non-local effects discussed in Sections 6 and 7. In this limit, the fractional framework must recover the known results of classical Resonant Relaxation (RR) theory, which serves as a consistency check for the model.
@@ 5.1 The Nature of the Stochastic Process
The limitation stems from the characterization of the stochastic process. Eq. (18) fundamentally describes Brownian motion or generalized Brownian motion (if noise is correlated but Gaussian), which relies on the Central Limit Theorem (CLT) and requires stochastic steps with finite variance. The physical description of RR adopted here corresponds to a Lévy flight (Eq. 8), a process where the variance of the step size is infinite (Section 2). The evolution of such a system converges to a Lévy stable distribution (Generalized Central Limit Theorem, GCLT), rather than a Gaussian distribution.
While treatments such as B. Bar-Or & T. Alexander (2014) correctly handle the temporal correlations of the noise, they explicitly assume the noise amplitude is Gaussian. In their formalism (Section 2.2 of B. Bar-Or & T. Alexander, 2014), the background potential terms ($h_{nm}^{l}$) are described as time-dependent Gaussian random variables by invoking the CLT, justified by the superposition of forces from $N\gg 1$ stars. As argued in Section 2, this justification is potentially flawed because the CLT requires finite variance of the individual contributions. If the variance of gravitational torques is infinite, the GCLT applies, leading to Lévy statistics, regardless of how large $N$ is. This assumption inherently maintains finite variance for the resulting stochastic process and necessitates a description via a local FP equation. This contrasts with the Lévy flight model where the GCLT applies.
Furthermore, one may argue that a pure power-law jump distribution (Eq. 8) is unphysical, as it allows for arbitrarily large jumps. A more physical model might involve “tempered” Lévy flights, where the power law is suppressed by an exponential factor, $P(\Delta J)\sim|\Delta J|^{-1-q}e^{-\lambda J}$ (M. M. Meerschaert & A. Sikorskii, 2011). This process, which converges to a “tempered stable” distribution, possesses finite variance. However, it is still fundamentally non-local and is described by a tempered space-fractional equation utilizing a tempered fractional operator $\mathbb{D}_{x}^{\alpha,\lambda}$ (M. M. Meerschaert & A. Sikorskii, 2011). Crucially, this operator remains an integro-differential operator, maintaining the non-local structure required to model the jumps. The local Fokker-Planck equation, which imposes a Gaussian propagator, is therefore not the correct description even for this more physically realistic, non-local process. This reinforces our central claim: a local differential operator is likely an inappropriate choice for describing Lévy-like dynamics, whether pure or tempered.
@@ 7.1 The Schwarzschild Barrier Hypothesis in the Adiabatic Invariance and Local Suppression
Classical two-body relaxation (TBR) is driven by stochastic encounters and does not depend on orbital coherence. It is entirely unaffected by the SB. However, TBR is much slower than RR in realistic nuclei.
The mechanism for RR suppression is understood as Adiabatic Invariance (AI). As detailed by B. Bar-Or & T. Alexander (2014), when the precession frequency exceeds the maximal variability frequency of the (assumed smooth and Gaussian) background noise, the effective diffusion coefficient is exponentially suppressed. This suppression was observed in N-body simulations using Post-Newtonian approximations (D. Merritt et al., 2011) and A. S. Hamers et al. (2014), and interpreted as a barrier. The standard view, based on these local models, is that the SB creates a significant bottleneck, as transport across the barrier region is limited primarily to the slow rate of two-body relaxation (B. Bar-Or & T. Alexander, 2016).